OpenStudy (anonymous):

Word Problem: A tourist travels 1500 miles using two planes. The second plane averages 50mph faster than the first plane. The tourist uses the slower plane for the first 500 miles and the faster plane for the next 1000 miles. The total flying time is 6.5 hours. What is the speed of the first plane???

OpenStudy (anonymous):



OpenStudy (anonymous):

\[500mile*\frac{ hr }{ slow mile } + 1000mile*\frac{ hr }{ slow+50mile } = 6.5hr\] \[500mile*\frac{ hr }{ X mile } + 1000mile*\frac{ hr }{ X+50mile } = 6.5hr\]

OpenStudy (anonymous):

\[\frac{ 500 }{ x } + \frac{ 1000 }{ x+50 } = 6.5\]

OpenStudy (anonymous):

make sense? ya no? x = slow plane speed = 200 mph

OpenStudy (anonymous):

500/200 = 2.5hr 1000/(200+50) = 4hr ... i'm going to bed... hope you can 'see it' It has mostly to do with units, using miles and hours and the ratios between them.